# Paper Copilot: The Artificial Intelligence and Machine Learning Community Should Adopt a More Transparent and Regulated Peer Review Process

### v1

@@ Abstract
The rapid growth of submissions to top-tier Artificial Intelligence (AI) and Machine Learning (ML) conferences has prompted many venues to transition from closed to open review platforms. Some have fully embraced open peer reviews, allowing public visibility throughout the process, while others adopt hybrid approaches, such as releasing reviews only after final decisions or keeping reviews private despite using open peer review systems. In this work, we analyze the strengths and limitations of these models, highlighting the growing community interest in transparent peer review. To support this discussion, we examine insights from Paper Copilot, a website launched two years ago to aggregate and analyze AI / ML conference data while engaging a global audience. The site has attracted over 200,000 early-career researchers, particularly those aged 18–34 from 177 countries, many of whom are actively engaged in the peer review process. Drawing on our findings, this position paper advocates for a more transparent, open, and well-regulated peer review aiming to foster greater community involvement and propel advancements in the field.
Machine Learning, ICML

### v2

@@ Abstract
The rapid growth of submissions to top-tier Artificial Intelligence (AI) and Machine Learning (ML) conferences has prompted many venues to transition from closed to open review platforms. Some have fully embraced open peer reviews, allowing public visibility throughout the process, while others adopt hybrid approaches, such as releasing reviews only after final decisions or keeping reviews private despite using open peer review systems. In this work, we analyze the strengths and limitations of these models, highlighting the growing community interest in transparent peer review. To support this discussion, we examine insights from Paper Copilot (https://papercopilot.com/), a website launched two years ago to aggregate and analyze AI / ML conference data while engaging a global audience. The site has attracted over 200,000 early-career researchers, particularly those aged 18–34 from 177 countries, many of whom are actively engaged in the peer review process. Drawing on our findings, this position paper advocates for a more transparent, open, and well-regulated peer review aiming to foster greater community involvement and propel advancements in the field.
Artificial Intelligence, Machine Learning, Community, Paper Copilot, Transparency
@@ 7 Conclusion
8 Impact Statement
We offer a timely reflection on peer review practices in the AI / ML community and present actionable insights derived from large-scale community analytics. By openly sharing peer review metrics and fostering transparency-focused dialogue, we aim to empower early-career researchers, encourage broader community participation, and help shape the conversation around more accountable and inclusive review systems. We hope this work contributes to the development of future peer review policies that prioritize openness, fairness, and global accessibility—ultimately supporting a more equitable and trustworthy scientific ecosystem in AI / ML research.
9 Acknowledgement
I would like to express my sincere gratitude to the global AI / ML research community for their trust, contributions, and engagement with Paper Copilot. Built entirely from scratch without institutional backing, this project was independently conceived, developed, and maintained during my spare time. I’m especially thankful to the many researchers who voluntarily shared data and insights, believing in the value of transparency and collective progress. I also thank my Ph.D. advisor, Prof. Yajie Zhao, for fostering an environment that supports academic freedom and self-directed exploration. While this work was carried out without direct involvement from my advisor or lab, her support allowed me the flexibility to pursue community-driven initiatives. All development, infrastructure, and operational costs were personally funded. More can be found at: https://papercopilot.com/acknowledgment/